OpenStudy (anonymous):

Ponderosa paint & Glass makes paint at 3 plants. It then ships the unmarked cans to a central warehouse. Plant A supplies 50% of the paint & past records indicate that the pain is incorrectly mixed 10% of the time. Plant B contributes 30%, with a defective rate of 5%. Plant C supplies 20%, with painted mixed incorrectly 20% of the time. If Ponderosa guarantees it's product & spent $10,000 replacing improperly mixed paint last year, how should the cost be distributed among the 3 plants?

OpenStudy (anonymous):

"the pain is incorrectly mixed" Oww.

OpenStudy (anonymous):

the paint is incorrectly mixed

OpenStudy (anonymous):

Let \(D\) be the event of mix error.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Let \(A, B, C\) be the probabilities they came from each respective plant, ok?

OpenStudy (anonymous):

Plant A supplies 50% of the paint \[ \Pr(A) =0.5 \]

OpenStudy (anonymous):

incorrectly mixed 10% of the time\[ \Pr(D|A)=0.1 \]

OpenStudy (anonymous):

This works for each plant the same way. Does this make sense?

OpenStudy (anonymous):

okay. yes make sense

OpenStudy (anonymous):

spent $10,000 replacing improperly mixed paint last year This basically means when we sort out this money, we're assuming mixing errors happened.

OpenStudy (anonymous):

So the amount given to plant A would be \[ \Pr(A|D) \]

OpenStudy (anonymous):

It works the same for the other plants by symmetry.

OpenStudy (anonymous):

So your first step is to find out what \[ \Pr(D) \]is.

OpenStudy (anonymous):

Can you do it? It's the same as the last question.

OpenStudy (anonymous):

okay i'll calculate it first

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

Okay really? I have to go soon. Please hurry.

OpenStudy (anonymous):

0.105 correct?

OpenStudy (anonymous):

Yes. Now do \[ \Pr(A|D) \]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[ \Pr(A|D) = \frac{\Pr(A\cap D)}{\Pr(D)}=\frac{\Pr(D|A)\times \Pr(A)}{\Pr(D)} \]

OpenStudy (anonymous):

You already calculated \(\Pr(D\cap A)=\Pr(A\cap D)\) when you tried to find \(\Pr(D)\)

OpenStudy (anonymous):

mm wait im confuse in P(A|D) given?

OpenStudy (anonymous):

oh wait sorry my bet

OpenStudy (anonymous):

thnks a lot!

OpenStudy (anonymous):

do i multiple all the answer with 10 000? @wio

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Notice that the probabilities all add up to 1. That is a good sign!

OpenStudy (anonymous):

yupp thank you!